# Generative Tools for Graphical Assets: Empirical Guidelines based on Game Designers’ and Developers’ Preferences

### v1

@@ 8.1 Theoretical and Practical Implications
As previously mentioned, whilst there is extensive work on the benefits and potentials of graphical asset generation tools in game design and development, there are limited studies on their use and fit within a design and development pipeline. The behaviour and preferences of game designers and developers has not been examined before in regards to how and when to use these tools. Hence, there are no existing guidelines on how to efficiently incorporate them into the design and development of a game.
@@
There are a number of practical implications that arise from the study results. GAG tools should be developed for incorporation in the early stages of design and development, these tools should be integrated into existing design and development environment as much as possible, they should priorities larger volumes of variations over quality, the users should be able to configure and have some control over the tools and the generated output artefacts, and finally the generated artefacts should comply with common data formats. These implications are derived from the behaviour and preferences of users as there is a strong preference towards using GAG tools in the early design stages for generating inspiration, exploring ideas and creating placeholders which are closer to the final intended artefacts. They prefer to have a large volume of variations which they can later manipulate and improve to create their chosen assets in their intended quality. They demonstrate a strong interest in being able to configure GAG tools, compliance of outputs with the common data formats and control over the functionality and output formats of the generated artefacts.
@@ 9 Concluding Discussions
This study has investigated the behaviour and preferences of game designers and developers towards the use of graphical asset generation tools, e.g. PCG, and other intelligent generative approaches. The study considers how these users interact with generative tools within the creation pipeline and takes into account the different contributing factors such as their experience, their role and the size of their respective teams. The results indicate that users prefer early design stage applications such as generating inspiration, exploring ideas and creating placeholders. As such, users prefer tools that can output larger volumes of artefacts at the cost of quality. No preference is shown for any one form of interface, instead, users express a need for smooth integration of GAG tools within their existing development environments. This includes complying with the design language of the host software (when integrated), presenting the necessary controls for the tool to function and outputting artefacts via common data formats. Users express a need for tool configurability, and find generation times of up to 10 minutes to be acceptable. These findings have led to a set of guidelines and recommendations for the use and integration of generative tools within game design and development pipelines, with considerations for the user’s intended application.
As the focus of this study was on interfaces and examining the pipelines, methods of interaction and expectations of users. It was not feasible to also fully implement the concepts presented by the UI, within the time-frame. This may have an affect on how the users perceive the tools, as they may already have biases toward what PCG tools are capable, or not capable of; especially regarding generation speed. To improve upon this, we suggest that future research observes these preferences on real generative tools. Furthermore, the testing of a generative tool that incorporates these findings would help to confirm these preferences.

### v2

@@ 9.1 Theoretical and Practical Implications
As GAG methods advance, their applications in games will continue to become more and more prevalent. Given our findings and resulting heuristics, the role of GAG tools in their present form within game pipelines is now clearer. With this, we adapt the existing GDLC framework (Ramadan and Widyani, 2013) to meet with the growing interest in GAG tools. We integrate these heuristics into the graphical asset GDLC, shown in figure 8, illustrating their relevance at different stages of the pipeline.
As previously mentioned, whilst there is extensive work on the benefits and potentials of GAG tools in game design and development (Li et al., 2020; Shen et al., 2020; Shen and Chen, 2021), there are limited studies on their use and fit within a design and development pipeline (Ling et al., 2024; Colado et al., 2023). While the use of creative tools is well established (Kasurinen et al., 2013; Seidel et al., 2019), the behaviour and preferences of game designers and developers with regard to GAG tools has not been explored. Hence, there are no existing heuristics or guidelines on how to effectively incorporate them into the design and development of a game.
@@
There are a number of practical implications that arise from the study results. GAG tools should be developed for incorporation in the early stages of design and development, these tools should be integrated into existing design and development environment as much as possible, they should prioritise larger volumes of variations over quality, the users should be able to configure and have some control over the tools and the generated output artefacts, and finally the generated artefacts should comply with common data formats. These implications are derived from the behaviour and preferences of users as there is a strong preference towards using GAG tools in the early design stages for generating inspiration, exploring ideas and creating placeholders. They prefer to have a large volume of variations which they can later manipulate and improve to create their chosen assets in their intended quality. They demonstrate a strong interest in being able to configure GAG tools, compliance of outputs with the common data formats and control over the functionality and output formats of the generated artefacts.
@@ 10 Concluding Discussions
This study has investigated the behaviour and preferences of game designers and developers towards the use of graphical asset generation tools, e.g. PCG, and other intelligent generative approaches. The study considers how these users interact with generative tools within the creation pipeline and takes into account the different contributing factors such as their experience, their role and the size of their respective teams. The results indicate that users prefer early design stage applications such as generating inspiration, exploring ideas and creating placeholders. As such, users prefer tools that can output larger volumes of artefacts at the cost of quality. No preference is shown for any one form of interface, instead, users express a need for smooth integration of GAG tools within their existing development environments. This includes complying with the design language of the host software (when integrated), presenting the necessary controls for the tool to function and outputting artefacts via common data formats. Users express a need for tool configurability, and find generation times of up to 10 minutes to be acceptable. These findings have led to a set of heuristics for the use and integration of generative tools within game design and development pipelines, with considerations for the user’s intended application.